OpenStudy (anonymous):

-12=6+9x this question pls

OpenStudy (anonymous):

x=-18/9 =-2

OpenStudy (anonymous):

is the right

OpenStudy (anonymous):

-12-6=9x -18=9x divide both sides by 9 -2=x